The two points(30,-25) and (50,65) determine a line
Found 2 solutions by checkley75, stanbon:
Answer by checkley75(3666)

THE SLOPE OF A LINE THROUGH THESE TWO POINTS IS
(Y2-Y2)/(X2-X1)
(65+25)/(50-30)
100/20
5 IS THE SLOPE (m)
NOW WE SUBSTITUTE ONE SET OF X & Y VALUES TO SOLVE FOR THR Y INTERCEPT(b) IN THE LINE EQUATION
Y=mX+b
-25=5*30+b
-25=150+b
B=-25-150
b=-175 THUS THE EQUATRION FOR THIS LINE IS
Y=5X-175

Answer by stanbon(75887)

The two points(30,-25) and (50,65) determine a line
slope = [65--25]/[50-30]= 85/20=17/4
Use the slope-intercept form to get the equation of the line.
y-65 = (17/4)(x-50)
y = (17/4)x -(17/4)50 +65
y = (17/4)x - 295/2
-------------------
Cheers,
Stan H.